Materials used by the Instrument Division of XPort Industries are currently purchased from outside suppliers at a cost of $327 per unit. However, the same materials are available from the Components Division. The Components Division has unused capacity and can produce the materials needed by the Instrument Division at a variable cost of $271 per unit.

a. If a transfer price of $298 per unit is established and 44,900 units of materials are transferred, with no reduction in the Components Division's current sales, how much would XPort Industries’ total income from operations increase?
$

b. How much would the Instrument Division’s income from operations increase?
$

c. How much would the Components Division's income from operations increase?
$
Answer #1

a. XPort Industries’ total income from operations increase by = ($327 - $271)*44900 = 2514400

b. Instrument Division’s income from operations increase by = (327-298)*44900 = $1302100

c. Components Division's income from operations increase by = (298-271)*44900 = $1212300